After their shopping trip mike spent 3/5 of his money.Eve spent 1/2 of hers and sue spent $50 they then had the same amount of money left.if the total amount lefr was $54 what was the total amount mike and sue had at first.
Answer by macston(5194)

At the end each had an equal part of $54, or $54/3=$18
Sue had spent $50, so she started with $18+$50=$68
Eve spent half, so she started with twice as much or 2($18)=$36
Mike spent 3/5, so he had 2/5 remaining, so he started with 5/2($18)=$45
CHECK:
Mike spent 3/5 and had $18 remaining
3/5($45)=135/5=$27 and $45-$27=$18
Given a number ($18) and its fraction of an original number (2/5) multiply by the reciprocal of the fraction (5/2) to obtain the original number. This is the same as dividing by (2/5)